# DreamO: A Unified Framework for Image Customization

### v3

@@ Abstract
Recently, extensive research on image customization (e.g., identity, subject, style, background, etc.) demonstrates strong customization capabilities in large-scale generative models. However, most approaches are designed for specific tasks, restricting their generalizability to combine different types of condition. Developing a unified framework for image customization remains an open challenge. In this paper, we present DreamO, an image customization framework designed to support a wide range of tasks while facilitating seamless integration of multiple conditions. Specifically, DreamO utilizes a diffusion transformer (DiT) framework to uniformly process input of different types. During training, we construct a large-scale training dataset that includes various customization tasks, and we introduce a feature routing constraint to facilitate the precise querying of relevant information from reference images. Additionally, we design a placeholder strategy that associates specific placeholders with conditions at particular positions, enabling control over the placement of conditions in the generated results. Moreover, we employ a progressive training strategy consisting of three stages: an initial stage focused on simple tasks with limited data to establish baseline consistency, a full-scale training stage to comprehensively enhance the customization capabilities, and a final quality alignment stage to correct quality biases introduced by low-quality data. Extensive experiments demonstrate that the proposed DreamO can effectively perform various image customization tasks with high quality and flexibly integrate different types of control conditions.
1 Introduction
@@
Due to the high-quality image generation and stable performance of diffusion models [14], substantial research efforts focus on controllable generation by leveraging their generative priors [34, 59]. Among these, image customization aims to ensure that generated outputs remain consistent with a reference image in specific attributes, such as identity [50, 53, 10], object appearance [18, 42, 12], virtual try-on [6, 29, 48], and style [55, 38, 52]. Despite the abundance of task-specific approaches, developing a unified framework for image customization remains a challenge.
@@
Addressing this challenge has long been a pursuit in the controllable generation community. For instance, early research Composer [16] jointly trains a diffusion model with multi-condition input, e.g., depth, color, sketch. Subsequently, adding extra control module/adapter [59, 34] based on pre-trained diffusion models greatly saves training costs. Following this idea, several UniControl [39, 60] methods propose to train additional control blocks to support general spatial control on the generation result. However, their control ability is restricted to some simple spatial conditions, and the interactions between conditions are rigid and have control redundancy. Recently, the DiT [36] framework greatly scales up the performance of diffusion models. Based on DiT, OminiControl [47] proposes to integrate image conditions by unified sequence with diffusion latent. It can perform various consistency preserving tasks, e.g., identity, color, and layout. Despite its advantages, OminiControl is trained separately on different tasks, struggling to process multiple conditions and model their interactions. For instance, it has difficulty with content insertion for multiple objects and virtual try-on with specified identities. Recently, OmniGen [54] trains a general generation control based on a pre-trained large language model [2] (LLMs). UniReal [3] achieves this through video generation pretraining followed by full-model post-training. Although these approaches benefit from the strong understanding ability to better interpret generation intentions, they require an enormous training cost involving hundreds of GPUs.
@@
In this paper, we design a unified image customization approach based on a pre-trained diffusion transformer (DiT). With a slight training cost, our method can support various consistency conditions (e.g., identity, subject, try-on, and style) and enables interactions among multiple control types, as shown in LABEL:fig:teaser. All capabilities are deployed on a single model. Specifically, we follow the unified sequence conditioning format introduced in OminiControl [47], and introduce a routing constraint on the internal representations of DiT during training. This ensures content fidelity and promotes the disentanglement of different control conditions. A placeholder strategy is also designed to enable control over the placement of conditions in the generated results. In addition, we construct large-scale training data covering multiple tasks and design a progressive training strategy. This enables the model to progressively acquire robust and generalized image customization capabilities.
@@
  • We introduce a progressive training strategy that facilitates convergence in multi-task and complex task settings. Additionally, we design a placeholder strategy to establish correspondence between textual descriptions and condition images.
@@
  • Extensive experiments demonstrate that our method not only achieves high-quality results across a broad range of image customization scenarios, but also exhibits strong flexibility in adapting to multi-condition scenarios.
2 Related Works
2.1 Diffusion Models
As a powerful generation paradigm, the diffusion model [14, 7] rapidly dominates the image generation community. Its high quality of generation and stable performance have been successfully applied to various tasks , e.g., text-to-image generation [44, 40, 35, 41] and image editing [13, 32, 56, 33]. To improve generation quality, some strategies are proposed, such as latent diffusion [41] and flow matching sampling [27]. Most early works utilize the UNet architecture as the diffusion backbone. Recently, the diffusion transformer [36] (DiT) architecture has emerged as a superior choice, offering improved performance through straightforward scalability. For instance, Stable Diffusion 3 [8] and Flux [1] utilize transformer architecture with unified and sequential text/image input.
2.2 Controllable Image Generation
Advancements in diffusion models drive the rapid development of controllable image generation. In this community, text is the most fundamental conditioning input, e.g., Stable Diffusion [41], Imagen [45], and DALL-E2 [40]. To achieve accurate spatial control, some methods, e.g., ControlNet [59] and T2I-Adapter [34], propose adding control modules on pre-trained diffusion models. UniControl [39, 60] propose to unify different spatial conditions with a joint condition input. In addition to spatial control, some unspatial conditions are also studied. The IP-Adapter [57] utilizes cross-attention to inject image prompt in the diffusion model to control some unspatial properties, e.g., identity and style. In addition to these representative works, other related attempts [9, 12, 15, 23, 25, 30, 31] also help to broaden the scope of controllable image generation.
Recent advancements in DiT-based diffusion models further promote the development of controllable image generation. For instance, in-context LoRA [17] and OminiControl [47] introduce a novel approach by concatenating all input tokens (i.e., text, image, and conditions) and training LoRA with task-specific datasets for various applications. Subsequently, OmniGen [54] and UniReal [3] optimize the entire diffusion model in multiple stages on larger-scale training data, achieving improved understanding of input conditions.
2.3 Cross-attention Routing in Diffusion Models
Existing studies (e.g., Prompt-to-Prompt [13]) demonstrate that text-visual cross-attention maps inherently establish spatial-semantic correspondence between linguistic tokens and visual generation. Notably, the routing pattern of cross-attention reveals a critical property: when computing attention scores between a specific condition feature and visual features, the response peaks consistently align with the spatial region of the corresponding subject in the synthesized image. Building upon this observation, UniPortrait [11] constrains the influence region of condition features for identity-specific generation in multi-face scenarios, and AnyStory [12] further extends this approach to subject-driven generation. In this paper, we investigate routing constraints within the DiT architecture.
3 Method
3.1 Preliminaries
The Diffusion Transformer (DiT) model [36] employs a transformer as the denoising network to refine diffusion latent. Specifically, in the input, the 2D image latent $\mathbf{z}_{t}\in\mathbb{R}^{c\times w\times h}$ is patchified into a sequence of 1D tokens $\mathbf{z}_{t}\in\mathbb{R}^{c\times(\frac{w}{p}\times\frac{h}{p})}$, where $(w,h)$ is spatial size, $c$ is the number of channels, and $p$ is the patch size. The image and text tokens are concatenated and processed by the DiT model in a unified manner. Apart from model architectures, more efficient sampling strategies (e.g., Flow Matching [27]) are also proposed. Unlike DDPM [14], Flow Matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the time step $t$, latent $\mathbf{z}_{t}$ is defined as: $\mathbf{z}_{t}=t\mathbf{z}_{0}+(1-t)\epsilon$, where $\mathbf{z}_{0}$ is the clean image, and $\epsilon\in\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised latent $\mathbf{z}_{t}$, timestep t, and condition $y$. The objective is to minimize the mean squared error loss:
$L_{diff}=E[||(\mathbf{z}_{0}-\epsilon)-\mathcal{V}_{\theta}(\mathbf{z}_{t},t,y)||^{2}],$ (1)
where $\mathcal{V}_{\theta}$ refers to the diffusion model. The DiT framework and Flow Matching are widely used in some recent diffusion models, such as Stable Diffusion3 [8] and Flux [1].
@@ 3.1 Preliminaries
Figure 1: Overview of our proposed DreamO, which can uniformly handle commonly used consistency-aware generation control.
3.2 Overview
An overview of our method is presented in Fig 1. Specifically, we utilize the Flux-1.0-dev [1] as the base model to build a unified framework for image customization, e.g., style, identity, subject appearance, and try-on. Given $n$ condition images $\mathbf{C}=\{\mathbf{C}_{1},...,\mathbf{C}_{n}\}$, we first reuse the VAE [22] of Flux to encode the condition image to the same latent space as noisy latent. Subsequently, the 2D latent representations are transformed into 1D token sequences via patchification, and all tokens are concatenated along the sequence dimension before being fed into Flux. To enable the model to incorporate conditional image inputs, we introduce a dedicated mapping layer at the input of Flux to encode the corresponding conditional image tokens. The position embedding (PE) of these conditional tokens is aligned with that of the noise latent using Rotary Position Embedding (RoPE) [46]. Inspired by non-overlapping position embedding in OminiControl [47], we extend these embeddings along the diagonal in a similar fashion. To distinguish conditional latent from noisy image latent in the latent space, a learnable condition embedding (CE) is added to the conditional latent at the input stage. Additionally, to support multi-condition scenarios, a learnable index embedding (IE) [3] is introduced and added to each conditional latent based on its associated condition index. Following the design of OminiControl [47], we integrate Low-Rank Adaptation (LoRA) [43] modules into Flux as trainable parameters, enabling the model to better adapt to conditional tasks.
Figure 2: Visualization of cross-attention maps in subject-driven image generation. The first row shows results from a model trained without routing constraints, while the second row presents results from a model trained with routing constraints.
Figure 3: Visualization of the placeholder effect.
3.3 Routing Constraint
Inspired by UniPortrait [11] and AnyStory [12], in this paper, we design routing constraint in the DiT architecture for general image customization tasks. As illustrated in Fig. 1, within the condition-guided framework, cross-attention exists between condition images and the generation result:
$\mathbf{M}=\frac{\mathbf{Q}_{cond,i}\mathbf{K}^{T}}{\sqrt{d}},$ (2)
where $\mathbf{Q}_{cond,i}\in\mathbb{R}^{l_{cond,i}\times c}$ refers to the condition tokens of the $i$-$th$ condition image. $\mathbf{K}\in\mathbb{R}^{l\times c}$ is the tokens of noisy image latent. The cross-attention map $\mathbf{M}\in\mathbb{R}^{l_{cond,i}\times l}$ is a dense similarity between the $i$-$th$ condition image and the generation result. To obtain the global response of the condition image across different locations of the generated output, we average the dense similarity matrix along the $l_{cond,i}$ dimension, resulting in a response map $\mathbf{M}\in\mathbb{R}^{l}$, representing the global similarity response of the condition image on the generated result.
To visualize the image-to-image cross-attention map in the DiT framework, we conduct a toy experiment. Specifically, we use the Subject200k [47] dataset to train the controllable framework in Fig. 1 on the subject-driven generation task. The first row of Fig 2 shows its generation result and inner cross-attention map. Here, we choose three layers (i.e., $2$, $8$, $14$) in the double block of Flux to visualize the cross-attention map. It can be observed that under the DiT framework, the image-to-image attention exhibits a spatial pattern, concerning around the target object, albeit with coarse granularity. To constrain the image-to-image attention focus on the specific subject, MSE loss is employed to optimize the attention within DiT across condition images and the generation result:
$L_{route}=\frac{1}{n_{c}\times n_{l}}\sum_{j=0}^{n_{l}-1}\sum_{i=0}^{n_{c}-1}||\mathbf{M}_{i}^{j}-\mathbf{M}_{target,i}||_{2}^{2},$ (3)
where $i$ and $j$ represents the condition index and layer index. $n_{c}$ and $n_{l}$ is the number of conditions and number of layers, respectively. $\mathbf{M}_{target}$ refers to the subject mask in the target image.
@@ 3.3 Routing Constraint
$L_{holder}=\frac{1}{n_{c}}\sum_{i=0}^{n_{c}-1}||Softmax(\mathbf{Q}_{cond,i}\times\mathbf{R}^{T})-\mathbf{B}_{i}||_{2}^{2},\ \mathbf{R}=[\mathbf{T}_{\mathbf{r}_{1}},\mathbf{T}_{\mathbf{r}_{2}},...,\mathbf{T}_{\mathbf{r}_{n_{c}}}],$ (4)
where $\mathbf{T}_{\mathbf{r}_{1}}$ refers to the feature of $[ref\#i]$. $\mathbf{B}_{i}$ is a binary matrix, where the value is $1$ when the placeholder matches the condition image, and 00 otherwise.

### v4

@@ Abstract.
Recently, extensive research on image customization (e.g., identity, subject, style, background, etc.) demonstrates strong customization capabilities in large-scale generative models. However, most approaches are designed for specific tasks, restricting their generalizability to combine different types of condition. Developing a unified framework for image customization remains an open challenge. In this paper, we present DreamO, an image customization framework designed to support a wide range of tasks while facilitating seamless integration of multiple conditions. Specifically, DreamO utilizes a diffusion transformer (DiT) framework to uniformly process input of different types. During training, we introduce a feature routing constraint to facilitate the precise querying of relevant information from reference images. Additionally, we design a placeholder strategy that associates specific placeholders with conditions at particular positions, enabling control over the placement of conditions in the generated results. Moreover, we employ a progressive training strategy to ensure smooth model convergence and correct the generation quality of the final output. Extensive experiments demonstrate that the proposed DreamO can effectively perform various image customization tasks with high quality and flexibly integrate different types of control conditions. Project page: https://mc-e.github.io/project/DreamO
Diffusion-models, Image customization
submissionid: 1976journal: TOGjournalyear: 2025copyright: acmlicensedconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763956isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Computer vision
1. Introduction
@@
Due to the high-quality image generation and stable performance of diffusion models (Ho et al., 2020), substantial research efforts focus on controllable generation by leveraging their generative priors (Mou et al., 2024; Zhang et al., 2023). Among these, image customization aims to ensure that generated outputs remain consistent with a reference image in specific attributes, such as identity (Wang et al., 2024a; Xiao et al., 2024b; Guo et al., 2024), object appearance (Huang et al., 2024a; Ruiz et al., 2023; He et al., 2025), virtual try-on (Choi et al., 2024; Luan et al., 2025; Wan et al., 2025), and style (Xing et al., 2024; Qi et al., 2024; Wu et al., 2021). Despite the abundance of task-specific approaches, developing a unified framework for image customization remains a challenge.
@@
Early research Composer (Huang et al., 2023) jointly trains a diffusion model with multi-condition input, e.g., depth, color, sketch. Some methods (Qin et al., 2023; Zhao et al., 2023) train additional control blocks (Zhang et al., 2023; Mou et al., 2024) to support general spatial control on the generation result, which greatly saves training costs. However, their control ability is restricted to some simple spatial conditions, and the interactions between conditions are rigid and have control redundancy. Recently, the DiT (Peebles and Xie, 2023) framework greatly scales up the performance of diffusion models. Based on DiT, OminiControl (Tan et al., 2024) proposes to integrate image conditions by unified sequence with diffusion latent. It can perform various customization tasks, e.g., identity, color, and layout. Despite its advantages, OminiControl is trained separately on different tasks, struggling to process multiple conditions. Recently, OmniGen (Xiao et al., 2024a) trains a general generation control based on a pre-trained large language model (Abdin et al., 2024) (LLMs). UniReal (Chen et al., 2024b) achieves this through video generation pretraining followed by full-model post-training. However, we argue that high-quality, multi-concept image customization cannot be achieved by merely leveraging the general capabilities of large language models like OmniGen or video models such as UniReal. Instead, it requires specialized architectural designs. Currently, the research community lacks an efficient and effective method specifically tailored for image customization under multi-concept and multi-conditional scenarios.
@@
In this paper, we design a unified image customization approach based on a pre-trained DiT. With a low training cost and a single model, our method can support various conditions (e.g., identity, subject, try-on, and style) and enables interactions among different kinds of condition, as shown in 1. Specifically, we follow the unified sequence conditioning format introduced in OminiControl (Tan et al., 2024), and introduce a routing constraint on the internal representations of DiT during training. This ensures content fidelity and promotes the disentanglement of different control conditions. A placeholder strategy is also designed to enable control over the placement of conditions in the generated results. In addition, we construct large-scale training data covering multiple tasks and design a progressive training strategy. This enables the model to progressively acquire robust and generalized image customization capabilities.
@@
  • We introduce a progressive training strategy to facilitate convergence in multi-task and complex task settings. Moreover, we design a placeholder strategy to establish correspondence between textual descriptions and condition images.
@@
  • Extensive experiments demonstrate that our method not only produces high-quality results in various image customization tasks, but also exhibits strong flexibility in multi-condition scenarios.
@@
Figure 2. Overview of our proposed DreamO, which can uniformly handle commonly used consistency-aware generation control.
2. Related Works
2.1. Diffusion Models
As a powerful generation paradigm, the diffusion model (Ho et al., 2020; Dhariwal and Nichol, 2021) rapidly dominates the image generation community. Its high generation quality and stable performance have been successfully applied to various tasks , e.g., text-to-image generation (Saharia et al., 2022a; Ramesh et al., 2022; Nichol et al., 2022; Rombach et al., 2022) and image editing (Hertz et al., 2022; Meng et al., 2021; Yang et al., 2023; Mou et al., 2023). Some strategies, such as latent diffusion (Rombach et al., 2022) and flow matching sampling (Lipman et al., 2022), are proposed to enhance the performance. Most early works (Rombach et al., 2022; Nichol et al., 2022) utilize the UNet architecture as the diffusion model. Recently, the diffusion transformer (Peebles and Xie, 2023) (DiT) architecture has emerged as a superior choice, offering improved performance through straightforward scalability.
2.2. Controllable Image Generation
Advancements in diffusion models drive the rapid development of controllable image generation. In this community, text is the most fundamental conditioning input, e.g., Stable Diffusion (Rombach et al., 2022), Imagen (Saharia et al., 2022b), and DALL-E2 (Ramesh et al., 2022). To achieve accurate spatial control, some methods, e.g., ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou et al., 2024), propose adding control modules on pre-trained diffusion models. UniControl (Qin et al., 2023; Zhao et al., 2023) propose to unify different spatial conditions with a joint condition input. In addition to spatial control, some unspatial conditions are also studied. The IP-Adapter (Ye et al., 2023) utilizes cross-attention to inject image prompt in the diffusion model to control some unspatial properties, e.g., identity and style. In addition to these representative works, other related attempts (Gal et al., 2022; He et al., 2025; Hua et al., 2023; Kumari et al., 2023; Li et al., 2023; Ma et al., 2024a, b) also help to broaden the scope of controllable image generation.
Recent advancements in DiT-based diffusion models further promote the development of controllable image generation. For instance, in-context LoRA (Huang et al., 2024b) and OminiControl (Tan et al., 2024) introduce a novel approach by concatenating all input tokens (i.e., text, image, and conditions) and training LoRA with task-specific datasets for various applications. Subsequently, OmniGen (Xiao et al., 2024a) and UniReal (Chen et al., 2024b) optimize the entire diffusion model in multiple stages on larger-scale training data, achieving improved understanding of input conditions.
2.3. Cross-attention Routing in Diffusion Models
Existing studies (e.g., Prompt-to-Prompt (Hertz et al., 2022)) demonstrate that text-visual cross-attention maps inherently establish spatial-semantic correspondence between linguistic tokens and visual generation, i.e., the similarity response aligns with the spatial region of the corresponding subject in the generated result. Building upon this observation, UniPortrait (He et al., 2024) constrains the influence region of condition features for identity-specific generation in multi-face scenarios, and AnyStory (He et al., 2025) further extends this approach to subject-driven generation. Some recent works (Yamaguchi and Yanai, 2024) show that the cross-attention map in the DiT framework also exhibits spatial properties. In this paper, we explore routing constraints in the DiT framework.
3. Method
3.1. Preliminaries
The Diffusion Transformer (DiT) model (Peebles and Xie, 2023) employs a transformer as the denoising network to refine diffusion latent. Specifically, in the input, the 2D image latent $\mathbf{z}_{t}\in\mathbb{R}^{c\times w\times h}$ is patchified into a sequence of 1D tokens $\mathbf{z}_{t}\in\mathbb{R}^{c\times(\frac{w}{p}\times\frac{h}{p})}$, where $(w,h)$ is spatial size, $c$ is the number of channels, and $p$ is the patch size. The image and text tokens are concatenated and processed by the DiT model in a unified manner. Apart from model architectures, more efficient sampling strategies (e.g., Flow Matching (Lipman et al., 2022)) are also proposed. Unlike DDPM (Ho et al., 2020), Flow Matching conducts the forward process by linearly interpolating between noise and data in a straight line. At the time step $t$, latent $\mathbf{z}_{t}$ is defined as: $\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\epsilon$, where $\mathbf{z}_{0}$ is the clean image, and $\epsilon\in\mathcal{N}(0,1)$ is the Gaussian noise. The model is trained to directly regress the target velocity given the noised latent $\mathbf{z}_{t}$, timestep t, and condition $y$. The objective is to minimize the mean squared error:
@@ 3.1. Preliminaries
where $\mathcal{V}_{\theta}$ refers to the diffusion model. The DiT framework and Flow Matching are widely used in some recent diffusion models, such as Stable Diffusion3 (Esser et al., 2024) and Flux (flu, 2023a).
3.2. Overview
An overview of our method is presented in Fig. 2. Specifically, we utilize the Flux-1.0-dev (flu, 2023a) as the base model to build a unified framework for image customization, e.g., style, identity, subject appearance, and try-on. Given $n$ condition images $\mathbf{C}=\{\mathbf{C}_{1},...,\mathbf{C}_{n}\}$, we first reuse the VAE (Kingma et al., 2013) of Flux to encode the condition image to the same latent space as noisy latent. Note that the size of the condition image is flexible. Higher resolutions are recommended for detail-rich images to preserve clarity, while lower resolutions are sufficient for images with fewer details, thereby reducing compression costs. Then, all tokens (i.e., image, text, condition) are concatenated along the sequence dimension and fed into Flux. To enable the model to incorporate the condition input, we introduce a condition mapping layer at the input of Flux. The position embedding (PE) of condition tokens is aligned with that of the noisy latent using Rotary Position Embedding (RoPE) (Su et al., 2024). Inspired by non-overlapping position embedding in OminiControl (Tan et al., 2024), we extend these embeddings along the diagonal in a similar fashion. In addition, we introduce a trainable and index-wise condition embedding (CE) $\mathbb{R}^{10\times c}$, which is directly added to condition tokens. Following OminiControl (Tan et al., 2024), we integrate Low-Rank Adaptation (LoRA) (Ryu, 2023) modules into Flux as trainable parameters.
Figure 3. Visualization of cross-attention maps in subject-driven image generation. The first row shows results from a model trained without routing constraints, while the second row presents results from a model trained with routing constraints.
3.3. Routing Constraint
Inspired by UniPortrait (He et al., 2024) and AnyStory (He et al., 2025), in this paper, we design routing constraint in the DiT framework for general image customization tasks. As illustrated in Fig. 2, within the condition-guided framework, cross-attention exists between condition images and the generation result:
(2) $\mathbf{M}=\frac{\mathbf{Q}_{cond,i}\mathbf{K}_{img}^{T}}{\sqrt{d}},$
where $\mathbf{Q}_{cond,i}\in\mathbb{R}^{l_{cond,i}\times c}$ refers to the condition tokens of the $i$-thth condition image. $\mathbf{K}_{img}\in\mathbb{R}^{l\times c}$ is the tokens of noisy image latent. The cross-attention map $\mathbf{M}\in\mathbb{R}^{l_{cond,i}\times l}$ is a dense similarity between the $i$-thth condition image and the generation result. To obtain the global response of the condition image in different locations of the generated output, we average the dense similarity matrix along the $l_{cond,i}$ dimension, resulting in a response map $\mathbf{M}\in\mathbb{R}^{l}$, representing the global similarity of the condition image on the generated result. To constrain the image-to-image attention focus on the specific subject, MSE loss is employed to optimize the attention within DiT across condition images and the generation result:
[…]
where $i$ and $j$ refer to the condition index and layer index. $n_{c}$ and $n_{l}$ is the number of conditions and number of layers, respectively. $\mathbf{M}_{target}$ refers to the subject mask for the target image. As shown in the second row of Fig. 3, after training with routing constraint, the attention of the condition image clearly focuses on the target subject, and the result shows improved consistency with the reference image in terms of details. In addition to improved consistency, this strategy also helps decoupling in multi-reference cases.
In addition to image-to-image routing constraint, we also design placeholder-to-image routing constraint to establish correspondences between textual descriptions and condition inputs. Specifically, for the $i$-thth condition, we append a placeholder [ref#i][ref\#i] after the corresponding instance name, e.g., “A women from [ref#1] and a woman from [ref#2] is walking in the park.”. During the training for multi-condition tasks, we calculate the similarity between the conditional image tokens and the placeholder tokens. The routing constraint ensures that the similarity between $\mathbf{C}_{i}$ and [ref#i][ref\#i] is $1$, while it is 0 for all other pairs:
[…]
where $\mathbf{K}_{text,i}$ refers to the text feature of [ref#i][ref\#i]. $\mathbf{B}_{i}$ is a binary matrix, where the value is $1$ when the placeholder matches the condition image, and 0 otherwise.